### v1

Abstract
Large Language Models (LLMs) have become integral to software engineering (SE), where they are increasingly used in development workflows. However, their widespread use raises concerns about the presence and propagation of toxic language—harmful or offensive content that can foster exclusionary environments. This paper provides a comprehensive review of recent research on toxicity detection and mitigation, focusing on both SE-specific and general-purpose datasets. We examine annotation and preprocessing techniques, assess detection methodologies, and evaluate mitigation strategies, particularly those leveraging LLMs. Additionally, we conduct an ablation study demonstrating the effectiveness of LLM-based rewriting for reducing toxicity. By synthesizing existing work and identifying open challenges, this review highlights key areas for future research to ensure the responsible deployment of LLMs in SE and beyond.
keywords:
Software Engineering, Large Language Models, Toxicity Detection, Toxicity Mitigation
1 Introduction
Large Language Models (LLMs), such as BERT, GPT-4, Claude, and LLaMA, have seen extensive applications across a wide range of domains, including education, research, healthcare, software engineering (SE), and finance [16]. For instance, LLM-powered chatbots and virtual assistants are transforming customer support by offering efficient and personalized services [38]. In the healthcare sector, LLMs are contributing to advancements in medical research, drug discovery, and patient care [7]. As LLMs become more integrated into society, their widespread use has drawn increasing attention to their societal implications, both positive and negative.
One pressing concern in the digital age is toxic language, which refers to harmful, offensive, or discriminatory language, including hate speech, personal attacks, threats, and harmful stereotypes. The prevalence of such toxic language in online platforms, such as social media and forums, creates hostile and exclusionary environments that hinder constructive dialogue and can cause significant harm to individuals [36]. Researchers are actively working to understand the underlying causes of toxic behavior and develop effective methods for its detection and mitigation. However, the field faces several challenges. Accurately identifying toxic language across diverse contexts and languages remains a significant obstacle. For example,  Radfar et al. [40] found that mildly offensive terms are more frequently used in tweets to express hostility between users with no social connection than between users who are mutual friends. Such complexities highlight the need for continued advancements in detection algorithms and comprehensive strategies to address toxicity.
LLMs, with their ability to process and analyze vast amounts of textual data, offer promising solutions for addressing these challenges [13, 34]. Unlike traditional methods, LLMs can detect nuanced patterns of harmful language, including subtle and context-dependent forms of toxicity, that might otherwise go unnoticed. Researchers have also proposed and implemented various strategies to detect and mitigate toxicity in LLM-generated content, underscoring the potential of LLMs to play a central role in promoting healthier online interactions.
This work synthesizes insights from a substantial body of publications on toxicity research conducted over the past four years. It delves into core topics such as toxicity detection and mitigation while also addressing critical auxiliary processes like corpus selection and data preprocessing. By consolidating findings from diverse studies, this paper provides a comprehensive overview of the current state of research and highlights gaps that require further exploration. One significant challenge with LLMs is their potential to generate toxic content, even in the absence of explicit prompts encouraging such behavior [20]. This issue stems from factors such as the presence of toxic content in training data [15], inherent algorithmic biases, or malicious exploitation. To better understand and address these issues, this paper proposes future research directions, including an ablation study that sheds light on the mechanisms underlying these behaviors and explores strategies for improvement.
Our main contributions include:
  • We systematically compare SE and non-SE datasets in the context of toxicity research
  • We categorize and summarize common annotation and preprocessing steps
  • We evaluate toxicity detection and mitigation solutions
  • We recommend using LLM-based rewriting approaches for toxicity mitigation, based on the results of our ablation study
The remainder of this paper is organized as follows: §2 introduces the use of LLM techniques in the SE domain, with a focus on toxicity analysis. §3 outlines common techniques for annotating, preprocessing, and utilizing datasets for toxicity detection and mitigation. §4 proposes improvements to pre-processing steps and toxicity detection. §5 details the setup, methodology, and results of our ablation study. The overall workflow is illustrated in Figure 1. Our code is available in an online repository: https://github.com/yangjack1998/SE-Mitigate
Figure 1: Overall Workflow of LLM Toxicity Detection and Mitigation
2 Background
This section presents the motivation and formal definition of toxicity studies within the SE community. Additionally, it highlights recent empirical works included in this review and outlines the criteria used for their selection.
2.1 Why is LLM in SE important
The integration of LLMs into SE has revolutionized various aspects of the field, offering substantial improvements in automation, efficiency, and decision-making. From code generation and summarization to commit classification and toxicity detection, LLMs are being increasingly leveraged to enhance software development workflows. This section explores the diverse applications of LLMs in SE, discussing how they contribute to key tasks such as classification and generation while also addressing their impact on developer interactions and the broader SE community.
2.1.1 Utilizing LLMs in SE
Software Engineering, a discipline that involves designing, developing, testing, and maintaining software, has hugely benefited from the advancement of LLMs. Many SE tasks can naturally be formulated as data analysis (learning) tasks such as classification tasks, regression tasks, and generative tasks [56]. These tasks are well-suited to being addressed by LLMs and hold significant potential for optimization and refinement through their application. For instance, automated code summarization, where the objective is to generate concise and descriptive natural language summaries for given code snippets or functions, can be naturally cast as a generative task.  Zhang et al. [59] propose a retrieval-based neural model to aid code summarization that improves the state-of-the-art methods by enhancing the model with the most similar code snippets retrieved from the training set.
Hou et al. [25] point out that among their collection of LLM studies for SE tasks, following generation tasks, which the majority (about 70.97%) of studies center around, there are classification tasks that take up 21.62%. For example, categorizing commits into maintenance tasks helps practitioners make informed decisions regarding resource allocation and effectively manage technical debt. Researchers continue to explore ways to enhance the accuracy of this task through the application of LLMs. A study [21] proposed leveraging BERT (Bidirectional Encoder Representations from Transformers) to classify commits into three categories of maintenance tasks: corrective, perfective, and adaptive. Their approach demonstrated an approximately 8% improvement in accuracy compared to the previous state-of-the-art model.
2.1.2 Benefiting SE community using LLMs
Toxic interaction can disrupt teamwork, discourage participation, and reduce productivity in software projects. Raman et al. [41] take an early step in implementing and illustrating a measurement instrument (an SVM classifier tailored for software engineering) to detect toxic discussions in GitHub issues. They also recognize the limited availability of large, software engineering-specific toxicity datasets, as well as the significant class imbalance they often exhibit (e.g., only 6 out of 1000 GitHub issues are toxic). To address this issue, Mishra and Chatterjee [34] explore the use of a zero-shot LLM (ChatGPT) for detecting toxicity in software-related text. Their findings show promising accuracy, underscoring the potential for integrating ChatGPT-enabled toxicity detection into developer communication channels. Currently, there is a lack of studies that go further to explore the mitigation of toxicity in software-related text using LLMs. To address this gap, a pilot study will be presented later in the paper to demonstrate the feasibility of such an approach.
In terms of the distribution of LLM usages in SE activities, among six phases of the Software Development Life Cycle (SDLC) (i.e., requirements engineering, software design, software development, software quality assurance, software maintenance, and software management), over half of the studies are observed in the software development domain [25]. Within the domain, code generation and code completion are two tasks of great interest. While LLMs’ assistance with these tasks could largely improve the efficiency of SE practitioners, they have more subtle implications that we should be aware of. An empirical study [20] finds that pretrained neural language models (LMs) are susceptible to producing racist, sexist, or other forms of toxic language, posing challenges to their safe deployment. The tendency of pretrained LMs to generate toxic text even from seemingly harmless prompts highlights the importance of improving data selection processes during pretraining. This concern extends to similar generative processes, including code generation, as they are fundamentally driven by LLMs trained for text generation tasks.
2.2 What is toxicity
Understanding toxicity is crucial for fostering healthy interactions in online and professional communities, including software engineering. While the term ”toxicity” is broadly used to describe harmful or disruptive language, its precise definition and measurement can vary across contexts. This section explores how toxicity is defined in existing research, highlights the challenges in creating consistent annotation guidelines, and discusses key metrics used to evaluate both human annotations and model performance in toxicity detection.
2.2.1 Definition
To study toxicity, defining the term is both essential and challenging, as it is broader and more subjective than other text classification problems, such as intent classification or spam detection [44]. The Perspective API [14] defines toxic content as ”texts that are rude, disrespectful, or unreasonable”. Gehman et al. [20] put more focus on the harm it creates, and describe toxic language as any form of speech or text that includes hate speech, offensive comments, harassment, or abusive behavior intended to harm, insult, or threaten others. A prior study [55] observed specific characteristics of toxicity definitions based on the domain of study. For example, in SE, the term is defined with more emphasis on its occurrence in conversations and distinctions such as ”offensive name-calling, insults, threats, personal attacks, flirtations, reference to sexual activities, and swearing or cursing” [46].
2.2.2 Metrics
Annotation guidelines play a crucial role in ensuring consistency in toxicity detection datasets. Annotators follow standardized instructions to label instances of toxicity, helping to train models with reliable data. For example, the models powering certain toxicity detection systems, such as the Perspective API, rely on manual annotations guided by published annotation guidelines [2]. However, while clear and well-defined guidelines improve annotation accuracy and consistency, they do not eliminate subjectivity. Biases in guideline design, cultural differences among annotators, and varying interpretations of toxicity can introduce inconsistencies that affect the generalization of the model in different contexts [18]. A critical challenge in toxicity research is determining whether guidelines strike the right balance between recall and precision: Overly broad definitions risk flagging benign language, while narrow definitions may fail to capture implicit toxicity.
To assess the reliability of these annotations, researchers measure inter-annotator agreement (IAA), a key metric that evaluates the consistency among annotators. Given the subjective nature of toxic language, ensuring high agreement is essential to produce reliable data sets [44, 46, 47]. High IAA indicates that annotation guidelines are effective and that annotators have a shared understanding of toxicity. Among the widely used methods for assessing IAA, Cohen’s $\kappa$ accounts for agreement occurring by chance and is calculated as $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$, where $P_{o}$ represents observed agreement and $P_{e}$ represents expected agreement by chance. Another commonly used metric, Krippendorff’s $\alpha$, generalizes inter-rater reliability by accommodating multiple annotators, diverse data types, and missing data. It is computed as $\alpha=1-\frac{D_{o}}{D_{e}}$, where $D_{o}$ represents observed disagreement and $D_{e}$ represents expected disagreement by chance. These measures are critical for validating the robustness and consistency of annotations in toxicity detection research. Yet, they have limitations—low agreement may indicate annotator disagreement, but it may also reveal deeper issues such as ambiguous labeling guidelines [42] or the need for more context-aware annotation strategies.
Beyond annotation reliability, evaluating model performance is essential for assessing the effectiveness of toxicity detection systems. Performance metrics such as precision, recall, accuracy, F1 score, and F0.5 score provide insights into different aspects of model behavior. These metrics ensure a comprehensive understanding of how well a model balances false positives and false negatives. Table 1 presents the confusion matrix, which serves as the foundation for calculating these metrics, while Table 2 defines and details their respective equations. By employing rigorous evaluation metrics, researchers can gauge the strengths and limitations of toxicity detection models and refine them for improved real-world application.
Table 1: Confusion Matrix
Predicted Positive Predicted Negative
Actual Positive True Positive (TP) False Negative (FN)
Actual Negative False Positive (FP) True Negative (TN)
Table 2: Model Performance Metrics
Metric Definition Equation
Precision Proportion of correctly predicted positive instances out of all instances predicted as positive $\frac{\text{TP}}{\text{TP}+\text{FP}}$
Recall Proportion of correctly predicted positive instances out of all actual positive instances $\frac{\text{TP}}{\text{TP}+\text{FN}}$
Accuracy Proportion of correctly predicted instances (both positive and negative) out of all instances $\frac{\text{TP}+\text{TN}}{\text{TP+FP+TN+FN}}$
F1 Score Harmonic mean of precision and recall, balancing the trade-off between the two metrics $2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
F0.5 Score Variation of F1 score that places more emphasis on precision than recall $(1+0.5^{2})\cdot\frac{\text{Precision}\cdot\text{Recall}}{0.5^{2}\cdot\text{Precision}+\text{Recall}}$
2.3 Empirical Corpus
Given the diverse nature of toxicity, ranging from offensive language to harmful behavior, datasets must be tailored to the specific type of content they aim to analyze. The purpose of this aggregation is to provide a comprehensive overview of the datasets used in toxicity research, with a focus on both SE and non-SE contexts. By distinguishing between SE and non-SE datasets, we can better understand the unique challenges and characteristics of toxicity in different environments. SE datasets, often derived from developer communication platforms such as code reviews, issue trackers, and forums, present unique challenges such as the need to account for technical jargon and collaborative discourse. On the other hand, non-SE datasets, which may include social media, news articles, or general-purpose comment sections, often feature broader forms of toxicity, including hate speech, cyberbullying, and harassment. This segmentation allows researchers to identify common trends, evaluate the applicability of models across domains, and highlight areas where further research and dataset development are needed.
Tables 4 and 5 provide a summary of key SE and non-SE datasets used in toxicity detection research, organized by their characteristics, sources, annotation methods, and evaluation metrics. This overview provides a foundational reference for researchers seeking to build on existing work, compare approaches, and identify gaps in dataset development to enhance toxicity detection systems.
2.4 Study Selection
To identify relevant articles, we conducted a systematic search across four major academic databases: IEEE Xplore, ACM Digital Library, Springer, and arXiv. Given our focus on recent advancements in toxicity research within software engineering, we restricted our search to articles published between 2020 and 2025. Our initial query included the keywords toxicity and software engineering to retrieve a broad yet relevant set of studies.
The selection process consisted of four sequential steps. First, we screened titles, abstracts, and keywords to ensure the studies were relevant, scientifically rigorous, and offered meaningful insights. Second, we assessed the credibility of each study by evaluating the publication venues. Third, we removed duplicate entries to maintain a unique dataset. Finally, we conducted a thorough qualitative assessment to identify the most relevant and influential studies. After selecting these core studies, we applied a snowballing approach, tracing citations both forward and backward, ultimately identifying a total of 63 papers for inclusion in this survey. The complete process is illustrated in Figure 2.
Figure 2: Study Identification and Selection Process
Table 3: Overview of Core Literature on Toxicity Study in Software Engineering
Index # Paper Title Publication Venue Publication Year Covered Domains Detection Mitigation
1 ToxiSpanSE: An explainable toxicity detection in code review comments ESEM  2023 SE datasets, Human annotator, Span-based detector, Domain-specific tailoring, LLM based model, Feature engineering Yes No
2 A benchmark study of the contemporary toxicity detectors on software engineering interactions APSEC  2020 SE datasets, Non-SE datasets, Human annotator, Algorithmic annotator, Domain-specific tailoring Yes No
3 Exploring ChatGPT for Toxicity Detection in GitHub ICSE  2024 SE datasets, LLM based model, Domain-specific tailoring, Human annotator Yes No
4 Stress and burnout in open source: Toward finding, understanding, and mitigating unhealthy interactions ICSE  2020 SE datasets, Domain-specific tailoring, Human annotator, Feature engineering Yes No
5 An Empirical Evaluation of the Zero-shot, Few-shot, and Traditional Fine-tuning Based Pretrained Language Models for Sentiment Analysis in Software Engineering IEEE Access  2024 SE datasets, Human annotator, Feature engineering, LLM based model, Domain-specific tailoring Yes No
6 Identification and mitigation of toxic communications among open source software developers ASE  2022a SE datasets, Human annotator, Data cleaning, Feature engineering, Domain-specific tailoring Yes No
7 Towards offensive language detection and reduction in four software engineering communities EASE  2021 SE Datasets, Human annotator, Feature engineering, Domain-specific tailoring Yes Yes
8 Realtoxicityprompts: Evaluating neural toxic degeneration in language models arXiv preprint  2020 Non-SE datasets, Algorithmic annotator, Data cleaning, LLM based model Yes No
9 Exploring ChatGPT for identifying sexism in the communication of software developers PETRA  2024 LLM based model, Prompt engineering Yes No
10 DeMod: A Holistic Tool with Explainable Detection and Personalized Modification for Toxicity Censorship arXiv preprint  2024 Non-SE Datasets, Human annotator, LLM based model, Domain-specific tailoring, Span-based detector Yes Yes
Table 4: Toxicity Related Datasets from SE community
Dataset Name Toxicity Task Data Source Source URL Size Annotation Paper Publication Venue Measurement
ToxiCR Dataset Toxicity Detection FOSS project code reviews and Gitter chat messages https://github.com/WSU-SEAL/ToxiCR 20k Manual labeling (2 of the authors)  2022a ASE Accuracy, Precision, Recall, F-score, Cohen’s $\kappa$
Benchmark Study Dataset Toxicity Detection FOSS project code reviews, Gitter chat messages, and Jigsaw test sample https://github.com/WSU-SEAL/toxicity-dataset/ 14k Manual labeling (2 of the authors)  2020 APSEC Accuracy, Precision, Recall, F-score, Cohen’s $\kappa$
ToxiSpanSE Toxic Span Detection ToxiCR Dataset https://github.com/WSU-SEAL/ToxiSpanSE 23k Manual labeling (2 independent annotators)  2023 ESEM Krippendorff’s $\alpha$, F1
GitHub Issues Dataset Toxicity Detection GitHub https://github.com/CMUSTRUDEL/toxicity-detector 2k Perspective API  2020 ICSE F0.5
Comment Dataset Toxicity Detection GitHub, Gitter, Slack, Stack Overflow https://archive.org/details/stackexchange 1,130k Perspective API and Regex  2021 EASE Precision
Table 5: Toxicity Related Datasets from Non-SE community
Dataset Name Toxicity Task Data Source Source URL Size Annotation Paper Publication Venue Measurement
Real Toxicity Prompts Toxicity Detection Reddit https://github.com/allenai/real-toxicity-prompts 22k Perspective API  2020 EMNLP Expected Maximum Toxicity and Toxicity Probability
Writing Prompts Toxicity Mitigation Reddit https://www.kaggle.com/datasets/ratthachat/writing-prompts 300k Toxicity detection API  2022 AAAI Toxicity Reduction
Offensive Language Identification Dataset (OLID) Offensive Language Detection Twitter https://sites.google.com/site/offensevalsharedtask/olid 14.2k 3 annotators per rater group per comment  2019 NAACL Per-class Precision, Recall, F1, Weighted Average
Hate Speech Dataset Hate Speech, Offensive Language Detection Twitter https://github.com/BenchengW/Offensive-Language-Detection-DL 25k 3 to 9 annotators per post  2020 AI2SD Accuracy
3 Methodology
This section outlines common methodologies for addressing toxicity in SE. We first discuss annotation techniques involving human and algorithmic annotators. Next, we describe preprocessing methods, including data cleaning, feature engineering and sampling techniques. Finally, we discussed various models from multiple perspectives, including general versus domain-specific, model architecture and task types.
3.1 Annotation
Accurate toxicity annotation is essential for building high-quality datasets, as the subjective nature of toxicity often requires nuanced human interpretation. This section explores various annotation methods, including human annotation and algorithmic approaches, while evaluating their respective strengths and limitations.
3.1.1 Human annotator
The involvement of human annotators is indispensable for capturing context-specific nuances in toxic content. To enhance annotation quality, researchers employ several strategies. One widely adopted approach is the use of multiple annotators with diverse backgrounds, which helps mitigate individual biases and improve perspective diversity. For instance, Sarker et al. [47] paired annotators of different genders to label samples, increasing the likelihood of balanced viewpoint representation.
To ensure consistent labeling, annotators often follow a well-defined rubric that outlines toxicity criteria. These rubrics can be pre-established or developed specifically for a given research context. For example, in an effort to tailor annotation guidelines to the SE domain, Sarker et al. [46] had a group of annotators analyze 1000 SE texts and empirically develop a rubric for identifying toxicity. This process not only clarified labeling standards but also reduced subjectivity by aligning annotators with detailed, research-specific rules. The resulting rubric was later reused in subsequent work [44], demonstrating its reliability and applicability across studies.
Beyond following structured guidelines, researchers also assess annotation quality through inter-annotator agreement metrics such as Cohen’s $\kappa$ or Krippendorff’s $\alpha$. The choice of metric depends on the complexity of the task and the specific research needs. For instance, Sarker et al. [47] utilized Krippendorff’s $\alpha$, which accounts for multiple annotators, missing values, and partial agreements—making it particularly suitable for nuanced tasks like toxicity annotation. By quantifying consistency between annotators, these metrics help validate dataset reliability. When disagreements arise, conflict resolution mechanisms are employed to establish a reliable ground truth. Typically, a third annotator or a consensus process determines the final label, ensuring that conflicting perspectives are reconciled.
Additionally, multi-label annotation offers a more nuanced alternative to traditional binary classification, which often oversimplifies the complexity of toxic language. Instead of strictly categorizing content as offensive or non-offensive, multi-label annotation captures contextual dependencies and annotators’ varying perspectives.  Arhin et al. [5] introduced an approach incorporating three label types—normal, offensive, and undecided—along with three labeling schemes (strict, relaxed, and inferred group labels). This method enables a more comprehensive and context-aware annotation process, ultimately leading to more robust toxicity detection models.
3.1.2 Algorithmic Annotator
To reduce the costs and time associated with manual annotation, algorithmic tools are increasingly employed in toxicity research. These tools enable efficient labeling of large-scale corpora, providing automated assessments that assist researchers in filtering and prioritizing data for analysis.
One widely used tool in this domain is the Google Perspective API [14], which assigns toxicity scores to text based on linguistic attributes such as ”toxicity,” ”insult,” and ”identity attack.” Developed by Google Jigsaw, this API leverages machine learning models trained on extensive datasets to identify harmful or offensive language patterns in online communication. The Perspective API is particularly useful for pre-labeling datasets, allowing researchers to streamline manual annotation efforts while gaining a granular understanding of toxicity intensity. Its scoring system facilitates the evaluation of language models and informs the development of mitigation strategies. Given its effectiveness, studies [13, 20, 41] frequently use Perspective API as a benchmarking tool for validating toxicity detection methods. Additionally, it is often integrated into hybrid annotation pipelines that combine algorithmic and human inputs, thereby improving data quality and model robustness.
However, despite its widespread adoption, the Perspective API has notable limitations and potential biases that impact its reliability and fairness. Studies have highlighted systematic biases in its scoring, particularly against marginalized groups. For example, Sap et al. [43] demonstrated that the API disproportionately flags African American Vernacular English (AAVE) as more toxic than standard English, underscoring the risks of algorithmic discrimination in automated moderation tools. Furthermore, the API’s lack of transparency in its training data and model updates presents challenges for reproducibility and fairness auditing. Bender et al. [9] emphasize that black-box models like Perspective API hinder researchers’ ability to critically assess how biases emerge, making it difficult to address or mitigate them. Without open access to the underlying data, researchers must rely on empirical testing to infer weaknesses, which limits their ability to propose targeted improvements.
In addition to machine learning-based approaches, researchers also employ simpler rule-based methods such as regular expression (regex) matching to detect explicit toxic content. Regex allows for the definition of patterns that capture specific linguistic constructs, including slurs, profanities, or abusive language, enabling rapid identification of harmful text. For instance, regex can be used to flag messages containing offensive words, including obfuscated spellings such as ”f**k.” The primary advantage of regex-based methods lies in their speed and cost-effectiveness, as they require minimal computational resources compared to more sophisticated machine learning models or APIs. However, regex-based approaches have significant limitations: they struggle to interpret context, irony, or subtle variations in language, often leading to high false positive or negative rates. Nozza et al. [37] discussed the complexities introduced by various obfuscation methods, such as character substitutions, which complicate the detection process and require regex patterns to account for numerous variations. For example, crafting regex patterns to identify variations like ”f**k” requires anticipating and encoding potential character substitutions. Despite these drawbacks, regex remains a valuable tool for filtering explicit toxicity, particularly when used in conjunction with more advanced annotation techniques.
3.2 Preprocessing
Preprocessing plays a pivotal role in reducing noise and standardizing inputs. A recent study [49] investigates the impact of popular preprocessing methods on transformer models. The findings indicate that, in some cases, with an appropriate preprocessing strategy, even a simple Naïve Bayes classifier can outperform the best-performing transformer by 2% in accuracy. Other studies provide direct evidence that certain preprocessing techniques enhance the effectiveness of hate and offensive speech detection [22] as well as toxicity detection [35].
3.2.1 Data Cleaning
Effective data cleaning is essential for preprocessing text in toxicity detection tasks, as extraneous and misleading content can interfere with model performance. One common step is URL removal, as URLs often introduce external references that may skew predictions. URLs are typically detected using regex patterns that match common formats such as ”http://” or domain names (e.g., www.example.com). Removing these links prevents models from relying on unrelated external content, ensuring they focus solely on textual features. This technique has been widely applied in prior studies [26, 6] to reduce biases. However, its overall effectiveness remains a topic of discussion, as studies have shown that URL removal has a negligible impact on sentiment classification accuracy and is often outperformed by more advanced preprocessing techniques [27].
Another important step is contraction expansion, where informal contractions like ”can’t” or ”won’t” are expanded into their full forms (”cannot” and ”will not”). This transformation standardizes text, making it easier for models to parse words correctly and recognize them as distinct entities. Studies such as Jianqiang and Xiaolin [27] have shown that contraction expansion improves accuracy and F1-scores in sentiment classification, demonstrating its relevance in text preprocessing pipelines. However, its effectiveness can vary across domains, as expanding contractions may introduce unnecessary verbosity or alter the tone of informal text, potentially impacting sentiment analysis in conversational contexts [1].
Symbol removal is another preprocessing step that enhances text clarity. Symbols such as ”@”, ”#”, or excessive punctuation (e.g., ”!!!”) are common in social media text but do not necessarily contribute meaningful information for toxicity detection. Their presence can introduce noise and cause models to overfit to irrelevant patterns. Removing unnecessary symbols [3, 52] helps streamline the data and ensures that predictions are based on meaningful language rather than stylistic artifacts. Similarly, repetition elimination addresses exaggerated expressions commonly found in informal text, where characters (e.g., ”loooove” → ”love”) or words (e.g., ”really really good”) are repeated. If not addressed, models may treat these variations as distinct features, leading to unnecessary complexity. By simplifying repeated elements [8], we maintain the intended meaning while improving model consistency. However, it’s important to note that while symbol removal can reduce noise, it may also strip away context that is crucial for understanding the tone or intent behind a message. For instance, emoticons often convey contextual meaning, which can be valuable for sentiment analysis [1]. Thus, indiscriminate removal of symbols may risk losing important nuances in informal text, particularly on platforms where symbols play a significant role in communication.
For domain-specific datasets, particularly those containing programming-related text, additional preprocessing steps are necessary to distinguish between toxic language and structural code components. One such technique is identifier splitting, which breaks down compound variable names into separate tokens. Many programming identifiers, such as ”isToxic”, use camelCase or snake_case formatting, making them difficult for models trained on natural language text to interpret. Splitting these identifiers into distinct words (”is” and ”toxic”) allows models to focus on meaningful linguistic elements rather than code syntax, enhancing their ability to detect toxicity [29].
Another preprocessing method specific to code-related datasets is programming keyword removal. Programming languages contain syntax elements such as keywords (”if”, ”else”, ”return”), operators (”+”, ”=”, ”&&”), and delimiters (”{”, ”}”, ”;”), which are crucial for code execution but generally irrelevant for toxicity detection. By removing these elements, models can concentrate on meaningful text components, such as comments or string literals, where toxicity is more likely to be present. This approach has been validated in studies like HaCohen-Kerner et al. [23], which identified the removal of HTML objects and syntax elements as a key step in effective preprocessing. Filtering out programming-specific tokens helps models assess toxicity more accurately by preventing structural code artifacts from influencing the analysis.
3.2.2 Feature Engineering
Feature engineering plays a crucial role in improving model performance by incorporating customized attributes that enhance the detection of toxic content. One key technique is adversarial pattern identification, which focuses on detecting and mitigating adversarial inputs designed to bypass toxicity filters. These inputs often include obfuscated slurs or profanities crafted to evade automated detection. For instance, Hartvigsen et al. [24] employ adversarial pattern generation to identify implicit toxicity, improving the robustness of toxicity classifiers against evasive language tactics. However, this approach can inadvertently perpetuate bias by focusing on specific, often marginalized, groups targeted by toxic language. It risks amplifying the very patterns it aims to detect, rather than addressing the root causes of toxic discourse. For example, a study [43] demonstrates that toxicity detection models trained on existing datasets often exhibit significant disparities in performance across different demographic groups, with higher false positive rates for language associated with minority identities. This indicates that focusing narrowly on adversarial pattern identification, without considering broader social context and potential biases, can exacerbate existing inequities in toxicity detection.
Another widely used approach is profane word counting, which involves calculating the density of harmful or offensive words in a given comment. This method provides a numerical representation of potential toxicity, serving as an important feature for classifier training. Research by Bhat et al. [10] outlines a taxonomy for workplace toxicity and highlights profanity as a key indicator of offensive language. Similarly, d’Sa et al. [17] include profane word counting in their preprocessing pipeline to enhance classification accuracy.
Embedding features leverage word or sentence embeddings, such as Word2Vec, GloVe [60], and BERT [12], to represent text in a high-dimensional continuous space. These embeddings capture semantic relationships between words, allowing models to better understand nuanced language patterns associated with toxicity. By transforming text into a structured numerical format, embedding features enhance the model’s ability to detect implicit and context-dependent toxic language. Embedding features, while powerful for capturing semantic relationships, risk amplifying societal biases present in their training data. Studies like Blodgett et al. [11] reveal that such biases can lead to disproportionate toxicity flagging of language associated with marginalized groups, necessitating careful mitigation strategies.
Finally, contextual features are crucial for analyzing user interactions and conversational history to disambiguate toxicity. Isolated toxic statements may appear ambiguous without considering the surrounding discussion. For example, Pavlopoulos et al. [39] investigate the importance of conversational context in toxicity classification, demonstrating that models incorporating thread structure and user history outperform those relying solely on individual comments. By integrating contextual features, models can differentiate between harmful intent and benign language, improving overall classification accuracy. However, processing long conversational histories introduces a significant computational burden, hindering real-time applications due to the memory and processing demands of capturing long-range dependencies. Research into transformer efficiency [53] highlights the ongoing efforts to address this challenge.
3.2.3 Sampling techniques
To ensure a balanced representation of toxic and nontoxic samples, some studies [33, 46] employ stratified sampling based on external toxicity scores, such as those from Google’s Perspective API. This approach prevents bias by balancing the distribution of toxicity likelihoods across subsets and ensures that the model is exposed to a diverse range of samples, avoiding overfitting to one class. However, while stratified sampling aims for representativeness based on pre-defined criteria, it may not necessarily select the most informative samples for model training. As an alternative,  active learning is a technique that strategically selects the most valuable data points for labeling, focusing on those that will maximize the model’s learning with minimal labeled data. It can achieve comparable or even better performance with significantly fewer labeled samples by strategically selecting those that maximize model learning [19]. Therefore, active learning could be explored as an alternative or complementary technique to stratified sampling, especially when the cost of labeling is a concern.
3.3 Detection & Mitigation
The detection and mitigation of toxicity in SE communication have evolved significantly, transitioning from simple, keyword-based approaches to sophisticated, context-sensitive models. Current strategies employ both general-purpose and domain-specific language models, reflecting an essential balance between broad applicability and specialized performance. Recent advances highlight the necessity of tailoring detection systems to SE-specific language, given the domain’s unique terminologies and nuanced communication patterns. The following sections detail these distinctions and explore various architectural designs and task-specific approaches in toxicity management within the SE domain.
3.3.1 General vs. Domain-Specific Toxicity
Large pre-trained language models, such as GPT-3, GPT-4, and Gemini, demonstrate notable flexibility in toxicity detection tasks due to their zero-shot and few-shot capabilities. Mishra and Chatterjee [34] showcased how ChatGPT effectively identified toxic comments on GitHub without requiring domain-specific fine-tuning, leveraging its training on diverse datasets. Similarly, Sultana and Begum Kali [50] demonstrated ChatGPT’s success in detecting gender biases and stereotypes, such as maternal insults and subtle stereotyping, in SE communications. Generative pre-training enables these models to adapt to a range of tasks with minimal input. However, performance evaluations reveal limitations: Shafikuzzaman et al. [48] showed that while GPT-4o performs reasonably well in smaller SE datasets, it falls behind fine-tuned models like seBERT when applied to larger datasets such as GIT and ARC. This highlights a critical trade-off between flexibility and performance when domain-specific nuances are involved.
Focusing on the specific requirements of SE discussions, domain-specific models fine-tuned on SE datasets consistently outperform general-purpose models in both accuracy and contextual relevance. Sarker [45] emphasized the importance of pre-processing SE-specific language, noting that general-purpose models often misclassify technical terms such as ”kill” or ”junk” as toxic. The other benchmark study [46] demonstrated how retraining in SE-specific data significantly improves performance. For example, experiments that evaluated tools like DPCNN and BFS after retraining on SE datasets such as Code Review and Gitter Ethereum revealed considerable improvements in key metrics such as precision, recall, and F score. The accompanying Figure 3 and Figure 4 highlight these findings, with F scores for DPCNN and BFS reaching 0.406 and 0.366, respectively, on Code Review, and 0.652 and 0.660 on Gitter Ethereum. These metrics underscore the value of retraining models to handle SE-specific language more effectively.
In addition, Sarker et al. [47] introduced a span-based toxicity detector that identifies specific toxic phrases in the SE text, providing improved interpretability for moderators. Beyond binary classification, Cheriyan et al. [13] analyzed toxicity across platforms like GitHub and Stack Overflow, advocating for SE-specific context-aware solutions to reduce conflicts and enhance community dynamics. Recent benchmarks also highlight seBERT’s macro F1-scores exceeding 90% on datasets like GIT, reflecting its ability to capture domain-specific nuances with unparalleled precision [48]. These examples emphasize that models tailored to the SE domain are essential for achieving superior performance in detecting and mitigating toxicity.
Figure 3: Code Review Dataset (Comparing Methods Before & After Retrain on SE Dataset)
Figure 4: Gitter Ethereum Dataset (Comparing Methods Before & After Retrain on SE Dataset)
3.3.2 Architectural Difference
Traditional models offer a straightforward approach to toxicity detection in SE discussions but often fail to account for contextual nuances. Early SE toxicity detection often relied on lexicon-based approaches or classical machine learning algorithms like SVMs. Lexicon-based approaches use predefined dictionaries of words or phrases associated with specific sentiments or toxicity levels. For example, terms like ”kill” or ”junk” might be flagged as toxic based on such a dictionary, without considering their contextual meaning [28]. These methods, while easy to implement and understand, often misinterpret technical jargon or colloquial expressions common in SE communication, leading to a high rate of false positives [46]Raman et al. [41] demonstrated the use of an SVM tailored for GitHub issue comments, revealing its potential for identifying toxicity trends across communities, but its performance was limited by its reliance on simple feature sets. Moreover, empirical evaluations have shown that while lexicon-based models can achieve acceptable precision on datasets like Code Review and Gitter Ethereum, their recall remains poor when handling nuanced or domain-specific expressions [44, 46].
In contrast, LLM-based models capitalize on generative pre-training to develop richer contextual understanding. These methods utilize transformer-based architectures (BERT, RoBERTa, GPT-3, GPT-4), trained on large-scale, generative pre-training objectives. For example, generative pre-training involves training models to predict masked words or the next token in a sequence, enabling them to capture rich semantic relationships in text [30, 51]. ToxiSpanSE [47], a fine-tuned RoBERTa-based model, achieved an impressive F1-score of 0.88 in detecting toxic spans in SE datasets, significantly outperforming earlier lexicon-based methods. Similarly, Shafikuzzaman et al. [48] evaluated GPT-3.5 and GPT-4, demonstrating their strong performance in zero-shot and few-shot settings on SE datasets like ARC and GIT. Another innovative example is the DeMod tool [31], which incorporates explainable AI capabilities to enhance both detection and mitigation of toxicity. By leveraging ChatGPT’s generative pre-training, DeMod offers fine-grained explanations and personalized modification suggestions to address toxic content effectively.
3.3.3 Task Type
Most toxicity detection approaches focus on a binary classification task, labeling each comment as either toxic or non‑toxic. For instance, the five datesets from 4 are all examples of binary-classified datasets from the SE community. Most models built using these datasets yield an overall yes/no decision without identifying which parts of the text are problematic. Models built using these datasets typically yield an overall yes/no decision without pinpointing which parts of the text are problematic. In contrast, one notable model adopts span‑level detection methods that identify the exact segments containing toxic language, thereby increasing interpretability and transparency [47].
Recent work has also explored multiclass classification to capture different levels of toxicity within a single input, offering an alternative to the conventional binary approach. For instance, one study focusing on maternal insults and stereotyping classified texts into multiple categories—such as “contains maternal insult,” “contains slight maternal insult,” “hard to say,” or “does not contain maternal insult”—and reported higher accuracy compared to binary classification [50]. However, the decision boundaries of these multiclass models are not entirely transparent. Moreover, settings such as the model’s temperature can also impact the results, leading to variability even when using the same prompt.
Beyond detection, several studies have investigated generative mitigation strategies. For example, a Conflict Reduction System (CRS) has been proposed that not only detects offensive language but also generates rephrased alternatives to reduce hostility in communities such as GitHub, Gitter, Slack, and Stack Overflow [13]. Similarly, the DeMod tool leverages explainable AI to provide fine‑grained detection results along with personalized content modification suggestions, thereby helping users mitigate toxicity before posting [31]. Despite its innovative nature, this approach remains in its early stages and requires further empirical validation on diverse and representative datasets.
4 Future Directions
As research on toxicity detection evolves, several key areas emerge as promising directions for future exploration. Advancements in preprocessing techniques can enhance data quality while preserving critical contextual nuances, improving model performance. Similarly, refining detection methods by integrating external knowledge and conversational context may help address challenges in accurately interpreting toxicity. Additionally, improving model interpretability and transparency through explainable AI techniques can foster trust and usability in automated toxicity detection systems. This section discusses these potential advancements, outlining future opportunities in preprocessing, detection, and explainability to guide ongoing research in this domain.
4.1 Preprocessing
A potential avenue for future work lies in refining data cleaning techniques to better preserve nuanced semantic meanings in text. For instance, while standard normalization processes often reduce elongated words such as ”loooove” to their base form ”love,” this approach may overlook subtle variations in emotional intensity or intent, potentially leading to the loss of valuable contextual or toxicity-related information. Similarly, symbol removal practices might benefit from greater specificity; instead of blanket removal, future research could explore methods to distinguish between symbols that contribute to textual semantics and those that introduce noise. Developing more granular preprocessing strategies that account for these subtleties could enhance the ability of toxicity detection models to capture contextually rich and emotionally complex patterns, ultimately improving their accuracy and robustness.
4.2 Detection
A major challenge in toxicity detection is accurately interpreting context, as certain phrases may be toxic in one setting but neutral or even positive in another. Future work could explore incorporating external knowledge sources, conversation history, or user intent modeling to enhance contextual understanding. For example, leveraging transformer-based models with multi-turn dialogue context could help distinguish between sarcasm, humor, and genuinely harmful content. A recent research [4] demonstrates that human perception of toxicity is highly dependent on preceding dialogue, and integrating this context into detection models significantly improves accuracy, highlighting the importance of conversational context in toxicity detection. Zhu et al. [61] present a framework that integrates topic modeling with external knowledge bases, offering a potential direction for improving toxicity classification by leveraging similar techniques to understand user intent and conversational context. Future research could build upon these findings by combining multi-turn dialogue modeling, knowledge-aware transformers, and socio-linguistic analysis to develop more nuanced toxicity detection systems capable of differentiating between targeted toxicity and benign discussions.
Many toxicity detection models function as black boxes, making it challenging for users and moderators to understand why content is flagged. To improve transparency, future research should prioritize explainable AI (xAI) techniques, such as attention visualization and SHAP (SHapley Additive exPlanations) [32]. SHAP assigns contribution scores to individual features, helping to identify which words or phrases influence a model’s toxicity predictions. By quantifying these contributions, it enhances interpretability and allow researchers to assess whether model decisions align with human judgment. Recent models increasingly emphasize interpretability; for example, ToxiSpanSE [47] improves transparency by detecting toxic spans in code review comments and highlighting specific toxic segments. Additionally, human-in-the-loop approaches that enable users to contest or refine model decisions could foster trust and usability while mitigating unintended consequences, such as excessive or insufficient content flagging.
5 Ablation Study
To evaluate the specific contributions of different prompting strategies and model choices, we conducted an ablation study. The goal was to measure their individual impacts on toxicity detection and mitigation tasks within SE discussions. The following sections describe our experimental setups and findings.
5.1 Experiments
Two experiments will be presented below: one for toxicity detection and another for toxicity mitigation. The detection experiment assessed the performance of LLMs in classifying toxic GitHub issues, while the mitigation experiment focused on their capability to rewrite individual toxic comments.
5.1.1 Experiment Setup for Toxicity Detection
To empirically evaluate the feasibility of leveraging LLMs for toxicity detection in SE communities, we conducted experiments using a dataset of 194 GitHub issues. Each issue was annotated as either containing toxic content or not. This dataset was created and shared by Ramen. In the original study, the authors trained an SVM classifier and tested their model on the same dataset.
The GitHub issues were collected using the GitHub API and GHTorrent and were manually labeled by the original authors. We evaluated two LLMs — GPT-4o and Gemini-1.5-flash — under two prompting conditions. In the few-shot setting, we selected three randomly chosen toxic examples and three non-toxic examples from the original paper’s training set as prompts. In the zero-shot setting, only a textual definition of toxicity was provided without any examples. After obtaining each model’s predictions (“Yes” for Toxic and “No” for Non-Toxic), we computed classification metrics, including accuracy, precision, recall, and F1-score for both classes, as well as overall accuracy.
5.1.2 Experiment for Toxicity Mitigation
In the mitigation experiment, we continued utilizing the dataset from the same source. The key difference from the detection experiment is that, instead of entire issues, the input consisted of individual comments. This granularity enables a more precise evaluation of the mitigation effects [41]. As illustrated in Figure 5, we employed a structured approach to ensure model independence by utilizing three distinct models, preventing over-reliance on any single approach and enabling a more robust mitigation strategy. Our pipeline first invokes a few-shot GPT-4o, selected due to its exceptional performance in the previous detection experiment 5.1.1, to identify if a comment is toxic. Prior studies have shown that large-scale language models such as GPT-4 outperform traditional classifiers in zero-shot and few-shot settings, particularly in identifying offensive and harmful language while maintaining high recall rates [48]. The ChatGPT series of models has been widely recognized for its effectiveness in handling nuanced toxicity detection, especially in software engineering discussions [34]. Then, the comment will be sent to the Gemini-1.5-flash model to rewrite the text as it has demonstrated strong performance in text generation while preserving semantic consistency. Unlike rule-based text sanitization methods, generative models like Gemini-1.5-flash provide more contextually appropriate rephrasings, ensuring that the rewritten text remains meaningful within the software engineering domain while effectively reducing toxicity [50]. Finally, to evaluate the effectiveness of the rewrites, we measured the semantic similarity using Sentence-BERT, a model designed for capturing sentence-level meaning beyond surface-level wording [48]. Sentence-BERT was selected due to its superior ability to assess the semantic consistency of rewritten comments compared to classical cosine similarity approaches, making it particularly effective for our mitigation pipeline [41]. This approach provided a robust estimation of how well the mitigated comments preserved the intent of the initial submissions rather than merely matching lexical choices.
Figure 5: Mitigation Experiment Workflow
Figure 6: Contrived demonstration of inputs and outputs from our mitigation pipeline.
5.2 Results
This section summarizes the experimental outcomes for toxicity detection and mitigation tasks. We first present results for toxicity detection, highlighting model performance across different prompting methods, followed by detailed findings from our mitigation experiments.
5.2.1 Detection Experiment Results
The performance metrics for the toxicity detection task are summarized in Table 6. All configurations demonstrated strong overall accuracy, with both GPT-4o and Gemini achieving around 0.87 accuracy in the zero-shot setting. Notably, GPT-4o zero-shot exhibited balanced precision and recall for both classes, with scores around 0.86–0.87. In contrast, GPT-4o few-shot showed a trade-off: it achieved a higher Toxic recall (0.90) compared to zero-shot (0.86), but the improvement came at the expense of slightly lower overall accuracy (0.84) due to a decrease in Non-Toxic recall (0.79).
Gemini zero-shot performed well across both classes, achieving the highest Non-Toxic recall (0.92) and strong Toxic precision (0.90), indicating a more conservative but effective approach. However, Gemini few-shot exhibited a more aggressive stance towards detecting Toxic content, reaching the highest Toxic recall (0.97). This came at the expense of increased false positives, leading to a drop in overall accuracy (0.77) and a significantly lower Non-Toxic recall (0.58).
Compared to the previous study by Ramen, which used an SVM classifier on the same test set and achieved 0.75 precision and 0.35 recall for Toxic content, the LLM-based approaches significantly improved both precision and recall. This highlights the potential of LLMs in toxicity detection, offering a more balanced and effective classification of Toxic and Non-Toxic content. While few-shot prompting generally enhanced Toxic recall, it also led to an increase in false positives, suggesting the need for careful calibration to optimize performance across both classes.
Table 6: Evaluation of Prompting Methods in Toxicity Detection
Model & Prompt Accuracy Non-Toxic P Non-Toxic R Non-Toxic F1 Toxic P Toxic R Toxic F1
GPT-4o zero-shot 0.87 0.87 0.87 0.87 0.86 0.86 0.86
GPT-4o few-shot 0.84 0.90 0.79 0.84 0.80 0.90 0.85
Gemini zero-shot 0.87 0.84 0.92 0.88 0.90 0.81 0.85
Gemini few-shot 0.77 0.95 0.58 0.72 0.68 0.97 0.80
Figure 7: Evaluation of Prompting Methods in Toxicity Detection
5.2.2 Mitigation Experiment Results
All toxic comments in our sample were successfully mitigated with a single rewrite, achieving a Mitigation Success Rate of 100.00%. The average number of attempts remained at 1.00, indicating that minimal rewriting was typically sufficient to transform a toxic comment into an acceptable form. Fig. 6 showcases examples of toxic comments alongside their rewritten versions. Upon examination of these examples, it becomes evident that LLMs can effectively remove toxicity while preserving the original meaning of the comments. Furthermore, for cases with relatively low similarity scores, we observed that these instances often contained little substantive content beyond their toxic elements. This suggests that when a comment is predominantly toxic, the rewritten version necessarily diverges more from the original.
5.3 Next Steps
The results of our ablation study demonstrate the potential of LLMs for both toxicity detection and mitigation. While existing models such as GPT-4 and Gemini have shown strong performance, particularly in non-toxic comment classification and mitigation success rates, there are several directions for further research to improve accuracy and applicability.
5.3.1 Domain-Specific Model Training
One key area for improvement is training LLMs on software engineering discourse. Prior research shows general-purpose models struggle with domain-specific terminology and context, leading to misclassifications [45]. Fine-tuning models using labeled SE datasets could enhance their ability to detect nuanced toxicity in developer communications.
5.3.2 Optimized Prompt Engineering
Our results suggest that prompt design significantly impacts the effectiveness of LLMs in both detection and mitigation. Previous studies have demonstrated that structured prompts can improve precision and recall in toxicity detection, as prompt clarity, specificity, and framing affect model performance [34]. Future work should explore advanced prompt engineering techniques, such as few-shot learning and chain-of-thought prompting, to optimize model outputs and minimize false positives or false negatives.
5.3.3 Explainable Toxicity Detection
While current approaches achieve high accuracy, they often lack interpretability, which is crucial in SE settings where developers require actionable insights. Research has highlighted the importance of integrating explanation mechanisms into LLM-based toxicity classifiers, such as providing rationale generation or highlighting specific toxic phrases [34]. Future work should investigate explainability frameworks that enhance model trust and usability while ensuring transparency in classification decisions.
6 Conclusion
As LLMs become increasingly integrated across diverse sectors—ranging from education and healthcare to software engineering—the challenge of managing toxic language remains ever more pressing. In this meta-review, we have synthesized recent advances in toxicity detection and mitigation, drawing on a wide spectrum of studies that span dataset construction, annotation protocols, and pre-processing techniques. This work echoes our introduction by highlighting the dual-edged nature of LLMs: while they offer robust contextual understanding and generative capabilities that can uncover nuanced toxic expressions, they also risk propagating harmful content if not carefully managed.
Our findings indicate that the efficacy of toxicity detection is critically dependent on the quality of datasets and the consistency of annotation standards, with pre-processing and sampling strategies playing a pivotal role in shaping model performance. Conventional detection methods, although effective in identifying overt toxicity, often fall short in capturing subtle, context-dependent expressions—an area where LLMs demonstrate clear advantages. Additionally, generative rewriting approaches show promise in mitigating toxicity by reformulating harmful content while largely preserving the original meaning, thus fostering more constructive communication.
In summary, this paper not only provides a comprehensive overview of current methodologies but also underscores existing limitations and the need for further refinement. Future research should focus on domain-specific model tuning, advanced prompt engineering, and enhanced interpretability to build more transparent toxicity mitigation systems. Through these efforts, we hope to pave the way for safer and more inclusive online environments.